Find all relative extrema and classify them. Use the Second Derivatives Test. f(x, y) = x² + 2xy – 2y? – 10x
Answer #1

: faid = 23 +2ny -ey= 10% . 4x = 3x + SJ-10 = 0 ① 1 = 9x - 490 (0) 8x = 4y =) x = 24 put in ☺ 324DC to= 이 X = - J1 + 4 (3)(0)fxx = 6x tyy = – 4 fxy = 2 - textyy (f ny) ² - ((*) (-4)- 4 = – 4 [6x +1] Dat (5,5) = -4 (685+1) = -4 (31) to selence (s, t )